OpenStudy (anonymous):

Find an equation of a line in standard form that goes through (1,-3) and has a slope of 7.

OpenStudy (anonymous):

an equation would be y=7x-9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I know the standard Ax+By=c. I just don't know what to do with it

OpenStudy (anonymous):

The slope is already given so you just need to put it in point slope form. That is y-y1=m(x-x1) We know the slope is 7 so we put that in, and we have the point (1,-3) so we put that in. We now have y+3=7(x-1) Now we distribute 7 on the right side. Do you know how to do that?

OpenStudy (anonymous):

ok you know the slope is 7 and it goes through the pt (1,-3). now plug these into y=mx+b i made an error with the answer i told you before the equation is y=7x-10